A bowl contains 27 M&M candies; 9 are red, 6 are green, and 12 are blue. A piece of candy is chosen, and its color is noted. Suppose you took 3 pieces of candy without looking. Define the event E: Each color is represented. What is P(E)?
Round your answer to 3 decimal places.
1
Expert's answer
2011-10-20T08:28:16-0400
Let RGB& be the event that the first candy was red, the second was blue, and
the third was red.
Similarly, define the events RBG, GRB, RBG, BRG,
BGR
Then these events are inconistent each with other and& E& is their union,
so
P(E) = P(RGB) + P(RBG) + P(GRB) + P(RBG) + P(BRG) + P(BGR)

Notice
that

& P(RGB)= 9/27 * 6/26 * 12/25 = (9*6*12)/(27*26*25)
& P(RBG)=
9/27 * 12/26 * 6/25 = (9*6*12)/(27*26*25) = P(RGB)

Similarly, it can be
shown that each of the events RBG, RBG, GRB, RBG, BRG, BGR has the same
probability
(9*6*12)/(27*26*25)

Hence
P(E) = 6 *
(9*6*12)/(27*26*25)
& = 6*9*6*12/(27*26*25)
& =
6*6*12/(3*26*25)
& = 6*6*12/(3*2*13*25)
& = 6*12/(13*25)
& =
72/325
& = 0.2215385
& = 0.222